OpenStudy (aaronandyson):

A body is travelling with uniform acceleration and the final velocity is sqrt(180-7x) where x is the distance travelled by the body.Then the acceleration is?

OpenStudy (phi):

I think you would do \[ v= \frac{dx}{dt}= (180 -7x)^\frac{1}{2} \\ a= \frac{dv}{dt} = \frac{d^2x}{dt^2} = \frac{1}{2} (180-7x)^{-\frac{1}{2} }\cdot -7 \frac{dx}{dt} \] we can replace dx/dt to get \[ a = -3.5 \cdot (180-7x)^{-\frac{1}{2} } \cdot (180-7x)^{\frac{1}{2} } \\ a= -3.5\]

OpenStudy (mayankdevnani):

Alternative Way :- \[\large \bf v^2=u^2+2as\] \[\large \bf v=\sqrt{u^2+2as}\] and we have given \[\large \bf v=\sqrt{180-7x}\] compare them

OpenStudy (mayankdevnani):

we get, \[\large \bf u^2=180\] \[\large \bf 2ax=-7x\] \[\large \bf a=\frac{-7}{2}=-3.5m/s^2\]

OpenStudy (aaronandyson):

kya h be yeh?nahi samjha muhje

OpenStudy (mayankdevnani):

bhai,equation of 3rd motion likh

OpenStudy (aaronandyson):

v^2 = u^2 + 2aS

OpenStudy (mayankdevnani):

tu usko aise bhi likh sakta h \[\large \bf v=\sqrt{u^2+2as}\]

OpenStudy (mayankdevnani):

right?

OpenStudy (aaronandyson):

ohk so on comparing v^2 = u^2 + 2as with the given equation we get v^2 = 180-7s u^2 = 180 2as = -7s 2a = -7 s = -3.5ms^-2

OpenStudy (mayankdevnani):

yes